Eliminate theta from the equation
1.lambda cos2theta=cos(theta+a),lambda sin2theta=2sin(theta+a)

Grade:8

1 Answers

sonika p
6 years ago
squaring and adding both the equations, we get, 
lambda2 cos2theta = cos2(theta +a)
lambda2 sin2theta = 4sin2(theta +a)
adding both the equations, 
lambda2 (1) = 1+ 3sin2(theta +a)
sin2(theta +a) = (lambda2 -1)/3
sin(theta +a) = root((lambda2 -1)/3)
theta = sin-1(root((lambda2 -1)/3)) – a